### v2

# Symmetry-broken charge-ordered ground state in CsV3Sb5 Kagome metal

Abstract
The newly discovered family of non-magnetic Kagome metals AV3Sb5 (A=K,Rb,Cs) offers a unique platform for studying the interplay between a charge density wave transition with superconductivity, non-trivial topology, and spontaneous time-reversal symmetry. Despite characterizing the charge density wave phase is crucial to understand and model all these exotic properties, it remains unresolved. In this work, we use first-principles calculations of the free-energy, incorporating both ionic kinetic energy and anharmonic effects, to resolve the atomistic phase diagram of CsV3Sb5 and its charge ordering structure. Our results reveal a competition between metastable stacking orders of the reconstructed vanadium Kagome layers, which are energetically favored to form exclusively a triangular-hexagonal arrangement, allowing the possibility of competitive different domains and chiral order. Consistent with experimental observations, we find that the $2\times 2\times 2$ and $2\times 2\times 4$ modulations are nearly degenerate in free energy, and are abruptly melted into the high-symmetry hexagonal phase around 90 K. The transition is first-order but compatible with a measurable phonon-softening. Remarkably, even if the six-fold and inversion symmetry are intrinsically broken by the out-of-plane stacking, this does not lead to measurable anisotropy in the in-plane conductivity as suggested by measurements.
preprint: APS/123-QED
The Kagome lattice has been a focus of interest across numerous fields in physics since its introduction in 1951 [1]. Composed of corner-sharing triangles in a hexagonal cell, it exhibits intriguing inherent electronic properties, such as frustrated magnetism [2, 3], flat bands [4, 5, 6, 3], and symmetry-protected Dirac cones [7, 8, 9, 10], which lead to topologically non-trivial effects [11, 10]. Among Kagome materials, the recently discovered non-magnetic AV3Sb5 family (A=K,Rb,Cs) [12] is particularly unique as it hosts a plethora of exotic quantum properties such as electronic topology [13], superconductivity [14], and spontaneous time-reversal symmetry breaking [15], all within an still unresolved charge-density wave (CDW) phase. The lack of consensus regarding the CDW phase hinders the understanding and modeling of these exotic properties. Indeed, observations such as the chiral transport [16, 17] as well as the superconducting [18, 19] and time-reversal symmetry breaking [20, 21] mechanisms remain open questions. Solving the puzzle of the charge reconstruction is thus a central focus at the moment.
All members of the AV3Sb5 family undergo a CDW phase transition at temperatures around 90 K. The nature of the emerging charge-ordered phases remains debated, particularly regarding its modulation pattern and symmetry preservation. While experimental evidences suggest a $2\times 2\times 2$ modulation for potassium and rubidium compounds [22], contradictory claims exist about the out of plane stacking of CsV3Sb5. Although most studies support a $2\times 2\times 2$ CDW reconstruction [7, 23, 24, 25, 26, 27], reports of a $2\times 2\times 4$ modulation [28, 29], mixed domains of both [30, 31, 22, 32], or even transitions between them [30] exist. Additionally, the symmetry of the charge-reconstructed phase remains uncertain. In K- and Rb-based structures, a six-fold symmetry breaking $\pi$-shifted triangular-hexagonal (TrH) ordering is well established [22]. However, for CsV3Sb5, alternative stacking orders, such as mixed TrH and star-of-david (SoD)-like arrangements, seem to be necessary to interpret angle-resolved photoemission spectroscopy (ARPES) [23] and x-ray measurements [22]. In fact, the very existence of an intrinsic six-fold symmetry remains unclear. While scanning tunneling microscopy (STM) [7, 33] observes a six-fold rotational symmetry breaking at the surface, corroborated by 120°-rotated nematic domains in magneto-optical Kerr measurements [34], a recent electronic transport study on strain-free devices indicates that in-plane anisotropy is highly strain-sensitive, and vanishes in the absence of external perturbations [8]. This raises the question of whether the six-fold symmetry breaking is an intrinsic property of the CDW itself.
Figure 1: Structural and dynamical properties of CsV3Sb5. (a) Atomic structure of the pristine high-symmetry phase. (b) Visualization of the Brillouin zone of the high-symmetry phase and its high-symmetry points. (c) Harmonic and SSCHA Hessian phonon spectra of the high-symmetry phase calculated at 40, 60 and 80 K. Imaginary modes are depicted with negative values. The high symmetry wave-vectors calculated explicitly and later used for the Fourier interpolation are highlighted in green.(d),(e),(f) Atomic structure of the TrH, $\pi$-TrH and 4TrH phases. In the case of TrH and $\pi$-TrH, their respective primitive cells are highlighted. (g) Visualization of the out-of-plane shift of the vanadium TrH layers that distinguish the emerging low-symmetry metastable phases.
The lack of consensus on the nature of the charge-reconstructed phase has motivated theoretical works in order to explain the origin of the CDW transition. The lack of evidence of phonon softening [35, 27, 36, 26], along with the observed discontinuity in the lattice parameters at $T_{CDW}=94$ K [37], challenges the conventional soft-phonon driven transition so well-established in other CDW materials like transition-metal dichalcogenides (TMDs) [38, 39, 40] and supports a first-order phase transition scenario. However, ab initio harmonic phonons in the high-symmetry phase reveal dynamical instabilities that coincide with the experimentally observed in plane-doubling of the unit cell [41, 42], supporting the presence of soft-phonon physics. This also underlines how standard calculations, which restrict to electronic Born-Oppenheimer (BO) energies and its second-order derivatives the analysis of the thermodynamic and dynamic stability of different CDW phases, completely break down. A recent study has indeed demonstrated that fully considering the ionic kinetic energy and anharmonicity is necessary to understand why the high-symmetry phase is stable and becomes the thermodynamic ground state above $T_{CDW}$ [43].
In this work, making use of first principle calculations including anharmonicity and ionic fluctuations within the stochastic self-consistent harmonic approximation (SSCHA) [44, 45, 46, 47], we resolve the temperature dependent phase diagram of CsV3Sb5, unveiling its charge-ordered structure below $T_{CDW}$. We demonstrate that the TrH ordering of the Kagome vanadium plane is the only candidate for the low-symmetry CDW phase. Our results reveal a competition between the $2\times 2\times 2$ and $2\times 2\times 4$ stacking phases, explaining the ongoing debate in the community, as both structures are dynamically stable and almost degenerate in free energy. Additionally, we show that the observed phenomenology behind the six-fold symmetry breaking and band-splitting in the material is accurately described due to such stacking disorder. By studying the renormalization of the phonon spectral function and free-energy with temperature, we show that, even if the transition is of first-order character, it is compatible with the presence of soft phonons.
Above $T_{CDW}$ the primitive structure of CsV3Sb5 belongs to the P6/mmm (No. 191) space-group. It consists of a Cs hexagonal unit-cell, inside which a Sb hexagonal- and a V Kagome-layer are sandwiched between Sb honeycomb-layers (see Fig. 1(a)). As shown in Fig. 1(b), at the harmonic level, this phase exhibits lattice instabilities at the zone-border wave-vectors $\mathbf{q}_{M}$=(1/2,0,0) and $\mathbf{q}_{L}$=(1/2,0,1/2) [41, 42] (we express $\mathbf{q}$ vectors in units of the reciprocal space lattice vectors). This means that
[…]
both for the lowest energy modes at $\mathbf{q}_{L}$ and $\mathbf{q}_{M}$, where $\omega^{(0)}_{\lambda}(\mathbf{q})$ is the harmonic frequency of mode $\lambda$, $V(\mathbf{R})$ is the BO energy, and $Q_{\lambda}(\mathbf{q})$ the order-parameter related to the $\bm{\epsilon}_{\lambda}(\mathbf{q})$ polarization vector (see Supplementary Material). Harmonic imaginary phonon frequencies in a given structure thus signal that this structure is not a minimum of the BO energy surface, but that it lowers $V(\mathbf{R})$ along the lattice distortion described by the order parameter.
[…]
fully considering the BO potential as well as the ionic kinetic energy $K$ and entropy $S_{ion}$. The variational parameters are the centroid positions $\bm{\mathcal{R}}$, which determine at the end of the minimization the most probable ionic positions, and the auxiliary force constants $\bm{\Phi}$, related to the broadening of the ionic probability distribution function centered at the centroid positions. In analogy to the harmonic case (see Eq. (1)), an imaginary phonon frequency obtained from the Hessian of the SSCHA free energy,
[…]
signals that a structure is unstable towards the distortion described by a given normal mode. The $\omega_{\lambda}(\mathbf{q},T)$ phonon frequencies, usually called SSCHA Hessian phonons, are temperature-dependent, include non-perturbative anharmonicity, and can be understood as the position of the peaks of the phonon spectral function in the static limit [46, 48, 49].
The temperature-dependent anharmonic spectra obtained from the Hessian phonons is shown in Fig. 1(c) for the high-symmetry phase. In agreement with previous findings [43], at temperatures above 65 K this phase becomes dynamically stable thanks to ionic entropy, i.e. it becomes a local minimum of the free energy. The phonon instability at L is the last to disappear. In a second-order phase transition scenario, one could argue that the CDW phase must adopt a $2\times 2\times 2$ modulation below 65 K as a direct consequence of the phonon instability at L. However, given the contradictory findings regarding the out-of-plane modulation and the most probable first-order character of the transition, such an argument appears overly simplistic. Notably, the whole LM phonon branch becomes unstable at temperatures close to the transition. This is an indication that $2\times 2\times X$ modulated phases with lower free-energy than the high-symmetry phase may be present, which can not be ignored as possible charge-reconstructed phases. The possibility that any ordered phase has a lower free energy than the P6/mmm one above 65 K cannot be discarded, which would be the case in a first-order phase transition.
Solving the CDW phase diagram at low temperatures, therefore, requires to calculate the free energy of model CDW distorted phases as a function of temperature. The lowest-energy phonon branch in the LM high-symmetry line predominantly corresponds to an in-plane distortion of the vanadium Kagome layer. At the L and M high-symmetry points, the distortion is commensurate with a $2\times 2\times 2$ supercell, and the lowest energy modes transform under the $L_{2}^{-}$ and $M_{1}^{+}$ irreducible representations, respectively. To explore all possible emergent phases, we investigate the local-minima of the BO energy surface, $V(\mathbf{R})$, by relaxing independently distorted lattice configurations along the $L_{2}^{-}$ and $M_{1}^{+}$ directions. This process reveals seven local-minima of $V(\mathbf{R})$ (see Supplementary Material) with a lower BO energy than the high-symmetry phase, offering plausible models of the CDW reconstruction. Among them, both SoD and TrH arrangements of the vanadium layer are present. Remarkably, none of these models is dynamically stable at the harmonic level, further corroborating the necessity of including anharmonicity at a non-perturbative level to describe the CDW phase diagram of CsV3Sb5.
In a second step, we relax these model CDW phases considering the ionic kinetic energy and anharmonicity within the SSCHA at 0 K, without imposing any symmetry constraints. In order to broaden the analysis, we also consider a new configuration obtained by an arbitrary distortion at the mid-point of the LM branch. This structure, with a $2\times 2\times 4$ modulation, lacks any symmetry (P1 (No. 1) space-group). The energy landscape obtained with Eq. (2) simplifies to just three local minima. In fact, all charge ordering patterns apart from the TrH, including the SoD, are thermodynamically unstable and transition into a TrH ordering. Indeed, the TrH reconstruction is found in three different phases distinguished by the stacking-order (see Fig. 1(d-f)): the TrH, $\pi$-TrH, and 4TrH phases. The TrH phase belongs to the P6/mmm (No. 191) space-group given that the charge is reconstructed without any out of plane modulation, preserving the hexagonal symmetry, and it is energetically favored by 1.53 meV/f.u with respect to the high-symmetry phase (see Fig. 2(a)). The $\pi$-TrH phase belongs to the Fmmm (No. 69) space-group, and shows an ABABA stacking order commensurate in a $2\times 2\times 2$ supercell that breaks hexagonal symmetry. It is favored with respect to the high-symmetry phase by 2.25 meV/f.u. The 4TrH phase is practically degenerate in energy with the $\pi$-TrH phase, it belongs to the P1 (No. 1) space-group, and has lost all its symmetries due to its ABACA stacking, which is commensurate in a $2\times 2\times 4$ supercell. While the TrH phase is metastable, we can conclude that the $\pi$-TrH and the 4TrH phases are the ground state CDW phases of the system. External perturbations such as strain or impurities may influence the presence of one or the other, naturally explaining the controversial presence of $2\times 2\times 2$ and $2\times 2\times 4$ orders in experiments [7, 23, 24, 25, 26, 27, 28, 29, 30, 31, 22, 32, 30], and supporting the possibility of having disordered domains of different stacking. The ground state CDW phase has, thus, a clearly broken hexagonal symmetry, and the 4TrH case, as it has lost all symmetries, can also be labeled as chiral.
Figure 2: Energetics and structural evolution as a function of temperature of the TrH stacking orders. (a) Free energy as a function of temperature of the possible different CDW phases (continuous lines), compared to the phonon collapse of the high symmetry phase (dashed lines). The free energy is computed in the warming up process, while frequencies in the cooling. Projection of TrH and $\pi$-TrH CDW crystal structures as a function of temperature onto the order parameters of $L_{2}^{-}$ (b) and $M_{1}^{+}$ (c) modes. Considering that these modes are non-degenerate at a given wave vector, but the star is composed of three wave vectors, the order parameter is represented as $\mathbf{Q}_{L,M}=(a,b,c)_{L,M}$, where $a,b,c$ represent the projection onto each of the vectors in the star.
Figure 3: Fermiology and direction dependent conductivity of the low temperature CDW stable and metastable phases of CsV3Sb5. (a) Illustration of the Brillouin zone of the TrH phase. High-symmetry points of the primitive Brillouin zone are highlighted in color, while those of the pristine high-symmetry phase in black and gray. (b) Cut of the Fermi surface at $k_{z}=0$ plane of the TrH phase visualized in the Brillouin zone of the high-symmetry phase. The smaller Brillouin zone of the TrH phase is shadowed in color. (c) Direction dependent in-plane anisotropy of the electrical conductivity $\sigma$ for the TrH phase. The anisotropy is defined as $1-\sigma(\theta)/\sigma(\theta=0)$, and is represented as a function of an angle $\theta$ defined by the transport direction and the lattice-vector $\mathbf{a}$. (d-f) and (g-i) same as (a-c) for the $\pi$-TrH and 4TrH phases.
As temperature increases, the free energy difference of the CDW candidates with respect to the pristine phase is reduced, and at 89 K the latter phase becomes the ground state, in a perfect agreement with the experimentally observed $T_{CDW}=94$ K value [37]. Remarkably, the $\pi$-TrH ordering persists up to 105 K, where it transitions abruptly into the high-symmetry phase (Fig. 2 (b-c)). Within this temperature range, while the high-symmetry phase is thermodynamically favored, the $\pi$-TrH and the 4TrH phases remain metastable. Upon cooling from ambient conditions, a similar phenomena occurs. The high-symmetry phase is energetically favored until 89 K, after which the $\pi$-TrH and 4TrH phases become the ground state. However, down to temperatures around 65 K, the high-symmetry phase is dynamically stable given that the frequency of the $L_{2}^{-}$ mode is still positive. Thus, our calculations reproduce the first-order character of the CDW phase transition, in agreement with experiments [37]. Even if it will be difficult to realize it experimentally given the small energy differences, our results allow a hysteresis region of almost 40 K.
Despite our calculations show that the CDW ground state structures, the $\pi$-TrH and 4TrH phases, have a clear broken hexagonal symmetry, transport experiments show no intrinsic anisotropy in the electrical conductivity [8]. In order to solve this apparent contradiction, we study the fermiology of these phases and compare it with the symmetric metastable TrH. The TrH phase, as a consequence of its six-fold symmetry, has a Fermi surface with hexagonal symmetry (see Fig. 3(b)). In contrast, both the $\pi$-TrH and the 4TrH CDW ground states intrinsically break the six-fold symmetry due to the out-of-plane stacking. The $\pi$-TrH corresponds to a face centered orthorhombic structure, where two distinct M points (M1 and M2) are observed with a clearly different fermiology (see Fig. 3(e)): M2 shows a Fermi pocket absent at M1. In the case of the 4TrH phase, despite the absence of six-fold symmetry, the Brillouin zone is quasi-hexagonal and the Fermi surface at a first glance appears to respect six-fold symmetry, even if three different M points exist in the zone borders of the low-symmetry Brillouin zone (see Fig. 3(g)). By making use of semi-classical Boltzmann theory, we analyze the impact of the different fermiology on the anisotropy of the electrical conductivity. Figs. 3(c,f,i) show the angular dependence of the anisotropy. While the TrH phase shows symmetry-imposed isotropic conductivity, the $\pi$-TrH phase and the 4TrH phases do not. The $\pi$-TrH phase, shows a -1 % of anisotropy that peaks along the $\mathbf{a}+\mathbf{b}$ direction. This small effect, is further reduced by the out-of-plane disorder, as the 4TrH phase shows anisotropy below -0.2 %. These results demonstrate that the breaking of the six-fold symmetry is perfectly compatible with an isotropic conductivity in CsV3Sb5, consistent with experiments [8]. Moreover, the out-of-plane disorder is also consistent with the band splitting observed in ARPES measurements [23], which was before attributed to a possible TrH-SoD stacking that turns out clearly thermodynamically unstable in our calculations (see Supplementary Material).
Figure 4: Renormalization of the phonon-spectra of the $\pi$-TrH phase with temperature. (a) Phonon spectra computed by a Fourier interpolation from a $2\times 2\times 2$ $\mathbf{q}$-grid of the primitive lattice of the $\pi$-TrH phase. The high symmetry wave-vectors calculated explicitly and later used for the Fourier interpolation are highlighted in blue. (b) Atomic structure obtained from the distortion of the $T_{1}^{+}$ lattice-instability. (c) Top-view of the reconstruction of the TrH vanadium layer obtained from the distortion of the $T_{1}^{+}$ lattice-instability. The primitive unit-cell is highlighted. (d) Phononic spectral-function including phonon-phonon interactions at $\Gamma$ for the $\pi$-TrH phase, where the contribution of the $\Gamma_{1}^{+}$ mode is highlighted. (e) Same as (d) at $T$, highlighting the $T_{1}^{+}$ mode.
The fact that the CDW ground state structures are consistent with transport and ARPES data support their validity. Nevertheless, their dynamical stability needs to be confirmed by calculating its SSCHA Hessian phonons in a larger grid of their Brillouin zone. Such anharmonic temperature-dependent phonon spectra is plotted in Fig. 4(a) for the $\pi$-TrH phase. Due to computational limitations, the 4TrH phase could not be considered. Starting from 0 K, the structure remains dynamically stable up to temperatures close to the CDW transition. The absence of lattice instabilities confirms that the CDW persists up to $T_{CDW}$, and, hence, if an order-disorder transition occurs from the $\pi$-TrH to the 4TrH phase, it must be of first-order character, consistent with experimental reports [30]. Interestingly, a lattice-instability appears at approximately 80 K at the T point ($\mathbf{q}=(1/4,-1/4,1/2)$ in units of the reciprocal lattice vectors of the high-symmetry phase or (1,1/2,1/2) in units of the reciprocal lattice vectors of the $\pi$-TrH phase), rather than at $\Gamma$. This phonon-mode, which transforms under the irreducible representation $T_{1}^{+}$, is non-degenerate. The resulting distorted structure is commensurate in a $4\times 4\times 2$ supercell of the high-symmetry phase, and belongs to the Cmmm (No. 65) space-group (see Fig. 4(b,c)). Even if this instability could signal the presence of another phase between 80 and 89 K, which does not seem to be observed experimentally [37], it could be related to a second CDW phase with a larger in-plane order that has been recently observed under compression [50, 51]. In fact, the $T_{1}^{+}$ mode shows a strong renormalization with isotropic pressure in our calculations, with a tendency to become more unstable (see Supplementary Material), consistent with experiments.
By calculating the phonon spectral function with the dynamical extension of the SSCHA theory [46, 48, 49], we predict that the softening of the $T_{1}^{+}$ mode should be observable by, for instance, inelastic X-ray experiments. As shown in Fig. 4(e), this mode undergoes a significant renormalization before collapsing at 80 K, but it remains separated from the rest of the spectra, allowing its spectral peak to remain well-defined still at 75 K. We perform a similar analysis for the $\Gamma_{1}^{+}$ phonon mode, the one that emerges as the condensation of the $L_{2}^{-}$ mode of the high-symmetry phase. This mode exhibits a marked red-shift from 93 to 78 cm-1 between 0 and 75 K, accompanied by an increase in broadening, which makes it practically vanish from the spectrum at 75 K (see Fig. 4(d)). This behavior aligns well with previous Raman measurements, which report a strong renormalization of a $\Gamma_{1}^{+}$ mode at similar energies [52].
In summary, by performing first-principles calculations including the ionic kinetic energy, entropy, and anharmonicity, we resolve the phase diagram of CsV3Sb5 and its CDW phase. Our results yield a first-order CDW phase transition at 89 K, in good agreement with experimental results. The obtained CDW ground-state ordering is inherently three dimensional, with the nature of the stacking being hard to resolve. However, being consistent with transport and Raman data, the obtained $\pi$-TrH and 4TrH phases are realistic models of the CDW, and offer new avenues to the understanding of its still mysterious superconductivity and time-reversal symmetry breaking.
Acknowledgements.
This project was supported by the Spanish Ministerio de Ciencia e Innovación (grants PID2022-142008NB-I00 and PID2022-142861NA-I00). I.E. acknowledges the Department of Education, Universities and Research of the Eusko Jaurlaritza and the University of the Basque Country UPV/EHU (Grant No. IT1527-22). M.G.-A. was supported by Jane and Aatos Erkko Foundation, Keele Foundation, and Magnus Ehrnrooth Foundation as part of the SuperC collaboration. M.A. acknowledges a PhD scholarship from the Materials Physics Center. The author acknowledges the technical and human support provided by the DIPC Supercomputing Center.